OpenStudy (anonymous):

Help please. Can someone please show me how to work this out. I will give a medal. Let f(x) = x - 2 and g(x) = x2 - 7x - 9. Find f(g(-1)). -21 -3 3 21

OpenStudy (bibby):

g(-1) = (-1)^2 - (7*-1) -9 plug that into f(x) hint: (-1)^2 = 1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then what do I do?

OpenStudy (bibby):

There's nothing else to do. g(-1) = (-1)^2 - (7*-1) -9 g(-1) = 1- (-7) - 9 Simplify that and plug it into f(x)

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

I am sorry I am confused can you show me how to simply it so I can figure out the correct answer?

OpenStudy (anonymous):

I have never had a teacher that actually teaches math they just give you work to do and that is it.

OpenStudy (bibby):

Try not blaming the teacher. Just learn the math any way you can. Why can't you simplify "1- (-7) - 9"?